OpenStudy (anonymous):

-9 = 5/17 n... can u help me figure out which answer it is? a. n= -153/5 b. n= 5/153 c. n= -17/45 d. n= -45/17

OpenStudy (anonymous):

-153/5

OpenStudy (tyteen4a03):

Divide both sides by 5/17. If you don't have a calculator handy, remember that \[\div \frac{ a }{ b } = \times \frac {b} {a}.\] @arisauotani Please stop giving out answers and not explain them. It violates the OpenStudy Code of Conduct. Thank you.

OpenStudy (anonymous):